# Multi-Scale Spatial-Temporal Hypergraph Network with Lead-Lag Structures for Stock Time Series Forecasting

### v1

@@ 4.3 Ablation Studies
Table 2 illustrates the unique impact of each module and contrasts them with two hypergraph-based baseline models, STHAN-SR and ESTIMATE. We have the following observations: 1) When the multi-scale fusion module in Section 3.5 and the total multi-scale module (Remove the Total Multi-Scale module, including Decomposition in Section 3.2 and Fusion in Section 3.5, respectively) are removed, the performance of the model decreases in both cases. The performance drop is more severe when the entire module is removed, indicating that while decomposition helps, the internal fusion of multi-scale information is key to further enhancing performance. 2) Ablating the lead-lag module in Section 3.4, which limits the model to only intra-industry correlations, causes a notable performance decline. This highlights the importance of modeling inter-industry lead-lag relationships. Even without this module, Hermes still outperforms STHAN-SR and ESTIMATE which only consider internal industry correlations, demonstrating the superiority of Hermes’s core architecture over other hypergraph-based approaches. 3) Removing the skip-connection module in Section 3.6 also reduces performance, confirming its crucial role in facilitating information transmission and gradient flow, thereby improving the model’s feature extraction capabilities.
Ablation NASDAQ NYSE S&P500
Model Component IC ICIR Prec@N SR IC ICIR Prec@N SR IC ICIR Prec@N SR
w/o Multi-Scale Fusion 0.038 0.464 0.521 1.374 0.028 0.421 0.545 1.354 0.031 0.248 0.531 1.517
w/o Total Multi-Scale 0.030 0.422 0.506 0.975 0.023 0.347 0.527 1.199 0.032 0.213 0.503 1.075
w/o Lead-Lag 0.039 0.472 0.526 1.451 0.028 0.420 0.540 0.225 0.037 0.201 0.519 1.628
w/o Skip-Connection 0.041 0.441 0.536 1.655 0.029 0.345 0.539 1.449 0.043 0.252 0.522 1.653
STHAN-SR 0.039 0.451 0.543 1.416 0.029 0.344 0.542 1.228 0.037 0.227 0.549 1.533
ESTIMATE 0.037 0.444 0.539 1.307 0.030 0.327 0.536 1.115 0.035 0.241 0.553 1.547
Hermes (Ours) 0.044 0.538 0.535 2.161 0.032 0.466 0.549 1.655 0.050 0.334 0.539 2.247
Table 3: Efficiency analysis on public stock datasets, concerning the training time, inference time (measured in milliseconds per batch under the condition of a batch size being 1), and memory usage (quantified in GB) of different models.
Dataset NASDAQ NYSE S&P500
Property STHAN-SR ESTIMATE Hermes STHAN-SR ESTIMATE Hermes STHAN-SR ESTIMATE Hermes
Training Time 257 ms 1,267 ms 215 ms 384 ms 1,462 ms 282 ms 297 ms 1,317 ms 211 ms
Inference Time 52 ms 163 ms 42 ms 73 ms 241 ms 61 ms 32 ms 109 ms 25 ms
Memory Usage 2.15 GB 9.62 GB 0.98 GB 2.82 GB 10.25 GB 1.14 GB 1.23 GB 4.52 GB 0.54 GB
@@ 4.4 Efficiency Analysis
In order to ensure good applicability of the model in practical applications, the runtime and memory usage of the model are crucial. We conducted an efficiency comparison analysis of Hermes and two other hypergraph-based methods, STHAN-SR and ESTIMATE—see Table 3.
@@ 5 Conclusions
Stock time series forecasting plays a pivotal role in the modern economy, serving as a key tool for investors, financial institutions, and policymakers to make accurate decisions. However, existing forecasting methods often fall short in fully considering the lead-lag between industries and multi-scale information, resulting in limited predictive performance. To address these challenges, this study introduces the Hermes framework. By ingeniously fusing multi-scale analysis with lead-lag modules within a hypergraph network, the Hermes framework significantly boosts the performance of stock forecasting. Specifically, the hyperedge-based moving aggregation module incorporates a sliding window and utilizes dynamic temporal aggregation operations to consider lead-lag dependencies among industries. Meanwhile, the hyperedge-based multi-scale fusion module decomposes raw data at multiple scales, and then employs cross-scale, edge-to-edge message passing to integrate information from different scales while preserving the consistency of each scale. Experimental results on multiple real-world stock datasets show that Hermes outperforms existing state-of-the-art methods in both efficiency and accuracy.
@@ A.1.1 Datasets
To validate the performance of Hermes, we use three publicly available stock market datasets, with the dataset statistics provided in Table 4. These datasets provide a diverse set of stock time series, allowing us to comprehensively assess the model’s effectiveness across different financial scenarios. We evaluate our approach using three real-world datasets from the US stock market. These datasets contain relatively complete stock-industry relationships or Wiki-based company relations. The NASDAQ and NYSE datasets (Feng et al., 2019b) include EOD data from 01/02/2013 to 12/08/2017, filtered from the respective markets. Abnormal patterns and penny stocks were removed, while preserving their representative properties, with NASDAQ being more volatile and NYSE more stable. The S&P500 dataset (Huynh et al., 2023) gathers historical price data and industry information for the companies listed in the S&P500 index from the Yahoo Finance database, covering the period from 01/04/2016 to 05/25/2022.
@@
A.1.2 Implementation Details
For fair comparison, all samples are generated by moving a 16 time steps lookback window. All experiments with Hermes are conducted using PyTorch (Paszke et al., 2019) in Python 3.9, and are executed on a server featuring an Intel(R) Xeon(R) Platinum 8358 CPU and an NVIDIA Tesla-A800 GPU. We use ADAM (Kingma & Ba, 2015) with an initial learning rate of 5e - 3. The “Drop Last” issue is reported by several researchers (Qiu et al., 2025b; 2024). That is, in some previous works evaluating the model on test set with drop-last=True setting may cause additional errors related to test batch size (Qiu et al., 2025c). In our experiment, to ensure fair comparison in the future, we set the drop last to False for all baselines to avoid this issue.
A.1.3 Metrics
Previous studies use different evaluation metrics, making it challenging to perform a comprehensive comparison of various methods. To provide a thorough assessment of the methods performance, we use four of the most commonly applied and reliable metrics: IC and ICIR, which are rank-based evaluation metrics; Prec@N, which measures accuracy; and SR, which is return-based.
@@ A.1.3 Metrics
  • Precision@N (Prec@N): Prec@N evaluates the precision of the top N short-term profit predictions from the model. This way, we assess the capability of the techniques to support investment decisions. For example, when N is 10, and the labels of 6 among these top 10 predictions are positive, then the Prec@10 is 60%.
    $Prec@N=\frac{TP@N}{N},$ (27)
    where $TP$ is the number of true positive classifications. In this study, $N$ is set to 10.
  • Sharpe Ratio (SR): Sharpe ratio (Sharpe, 1994) measures the profitability of the investment method and take into account risk.
@@ A.1.4 Baselines
We conduct a comparative analysis between our proposed Hermes and a range of baseline methods, which include prominent approaches in both machine learning and deep learning methods for STSF. Given that existing research has demonstrated the limited effectiveness of traditional statistical methods (Fan & Shen, 2024; Li et al., 2024), we exclude them from our comparative evaluation to focus on more competitive approaches.
@@
  • LSTM (Hochreiter & Schmidhuber, 1997) refers to the standard LSTM model, which operates on sequential data, including closing prices and moving averages over 5, 10, 20, and 30 days, to generate a sequential embedding. A fully connected (FC) layer is then employed to predict the return ratio.
@@
  • ALSTM (Feng et al., 2019a) integrates adversarial training and stochastic simulation into an enhanced LSTM model, allowing it to more effectively capture market dynamics.
@@
  • RGCN (Li et al., 2021) introduces a model that captures both positive and negative correlations among stocks using a correlation matrix computed from historical market data. The LSTM mechanism in RGCN layers helps mitigate over-smoothing issues when predicting overnight stock price movements.
@@
  • GAT (Veličković et al., 2017) employs Graph Attention Networks (GAT) to aggregate stock embeddings, which are encoded by a GRU, on the stock graph. By combining the advantages of Graph Neural Networks (GNN) and attention mechanisms, it enhances performance in large-scale graph settings.
@@
  • RSR (Feng et al., 2019b) introduces a novel neural network-based framework called Relational Stock Ranking, which addresses the stock forecasting problem in a learning-to-rank approach. Additionally, it proposes a new component in neural network modeling, called Temporal Graph Convolution, which is designed to explicitly capture the domain knowledge of stock relationships in a time-sensitive manner.
@@
  • STHAN-SR (Sawhney et al., 2021) introduces a novel Spatio-Temporal Hypergraph Attention Network that models inter-stock relationships of different types and strengths as a hypergraph for stock ranking. It combines temporal Hawkes attention with spatial hypergraph convolutions through hypergraph attention to capture both the correlations in the movements of related stocks and the temporal evolution of their historical features.
@@
  • ESTIMATE (Huynh et al., 2023) integrates temporal generative filters within a memory-based shared parameter LSTM network, enhancing the model’s ability to learn temporal patterns for each stock. Furthermore, it introduces attention hypergraph convolutional layers that utilize the wavelet basis—a convolution approach that leverages the polynomial wavelet basis to streamline message passing and focus on localized convolutions.
@@
  • Linear (Zeng et al., 2023) employs a straightforward approach by utilizing only fully connected layers to predict the final price. This method does not incorporate any advanced architectures or temporal dependencies, relying solely on basic linear transformations to generate the forecasting. Despite its simplicity, it serves as a baseline model for comparison with more complex models.
@@
  • StockMixer (Fan & Shen, 2024) introduces a lightweight yet effective MLP-based architecture for stock price forecasting. The model incorporates three key components—indicator mixing, time mixing, and stock mixing—which work together to capture the intricate correlations within the stock data.
@@ B.1 Traditional Stock Time Series Forecasting
Early stock forecasting methods rely mainly on statistical learning techniques to capture patterns and relationships in time series (Ye et al., 2016; Ariyo et al., 2014; Gupta & Dhingra, 2012; Elton et al., 2009). One of the most widely used techniques is the Autoregressive Integrated Moving Average (ARIMA) (Ariyo et al., 2014) model, which stabilizes data by combining Autoregressive components, Moving Average, and differencing, thereby modeling time series. Another commonly used traditional method is the Hidden Markov Model (HMM) (Gupta & Dhingra, 2012), which is used to describe unobservable states in the market and can detect changes in market conditions, making it suitable for analyzing market transitions. Although these methods provide predictive capabilities, they exhibit significant limitations when dealing with nonlinear, dynamic changes, and sudden events in financial markets. As a market environment becomes more complex, traditional statistical methods increasingly fail to provide sufficient predictive accuracy and robustness.
With the rapid development in machine learning (Li et al., 2025a; Yang et al., 2024; Sun et al., 2024; Niu et al., 2025; Lu et al., 2024a), machine learning methods for stock time series forecasting have emerged (Alkhatib et al., 2013; Nugroho et al., 2014; Kim, 2003). For example, Support Vector Machines (SVMs) (Kim, 2003) map data to high-dimensional spaces using kernel functions and find optimal separating hyperplanes in those spaces, which allows them to capture complex relationships and dynamics in financial markets. However, SVMs are highly sensitive to parameter selection and often struggles to capture long-term dependencies. Overall, although machine learning methods show considerable potential in stock forecasting, they still rely on manual feature engineering and model design. Moreover, these traditional stock forecasting methods have not delved into the exploration of lead-lag relationships and multi-scale modeling.
@@ B.2 Deep Learning Stock Time Series Forecasting
Recently, deep learning has made impressive progress in natural language processing (Liang et al., 2024; Li et al., 2025b; Zhou et al., 2024), computer vision (Li et al., 2025d; Lu et al., 2023; 2024b; Zhou et al., 2025), and other aspects (Li et al., 2025c; Wu et al., 2024; Wang et al., 2025). Studies have shown that learned features may perform better than human-designed features (Liu et al., 2025c; Qiu et al., 2025a; Dai et al., 2024a). Therefore, many deep learning-based STSF methods have been proposed and have received substantial extensive attention. Some studies use recurrent neural networks (RNNs) (Nelson et al., 2017; Qin et al., 2017; Akita et al., 2016; Rahman et al., 2019) to capture the historical development in individual stock prices and predict their short-term trends. For example, Long Short-Term Memory (LSTM) networks combined with historical stock data and text information are used to predict stock prices (Akita et al., 2016). In addition, due to the significant progress of Multi-Layer Perceptrons (MLPs) architecture in the general time series field, the DLinear (Zeng et al., 2023), CycleNet (Lin et al., 2025), and SparseTSF (Lin et al., 2024) show the benefits of this simple architecture for time series forecasting. Increasingly may proposals target stock forecasting performance by improving the MLPs architecture. For example, a series of studies (Fan & Shen, 2024; Lazcano et al., 2024; Tashakkori et al., 2024) enhance the learning ability of simple MLPs structures by utilizing the MLP-Mixer backbone, thereby improving forecasting accuracy. Next, the Transformer model has become increasingly popular due to the ability of its self-attention mechanism to capture long-term dependencies. For example, Master (Li et al., 2024) proposes a novel stock transformer for stock price forecasting to effectively capture stock correlation.
Some deep learning models have begun to explore lead-lag relationships and multi-scale modeling. For instance, StockMixer (Fan & Shen, 2024) engages in multi-scale modeling by segmenting the original time series into subsequence-level patches and mixing features at different scales. However, this approach lacks effective integration of information across scales; it merely concatenates information from different scales without thorough fusion. On the other hand, ESTIMATE (Huynh et al., 2023) takes into account lead-lag relationships through data-driven detection of leading and lagging time series clusters. Nonetheless, this method falls short in capturing fine-grained lead-lag dynamics, as clustering can only reveal global lead-lag relationships. Distinctive from the aforementioned approaches, Hermes employs a sliding window technique and utilizes dynamic temporal aggregation operations to consider fine-grained lead-lag dependencies among stock industries. Moreover, it features a multi-scale fusion module, which utilizes a cross-scale, egde-to-edge message-passing to integrate information from different scales while maintaining the consistency of each scale.

### v2

@@ 4.3 Ablation Studies
Table 2 illustrates the unique impact of each module. We have the following observations: 1) When the multi-scale fusion module in Section 3.5 and the total multi-scale module (Remove the Total Multi-Scale module, including Decomposition in Section 3.2 and Fusion in Section 3.5, respectively) are removed, the performance of the model decreases in both cases. This underscores the effectiveness of our multi-scale approach, confirming that both multi-scale decomposition and fusion are essential to the model’s predictive power. 2) The ablation of the lead–lag module in Section 3.4 causes a substantial decline in performance, demonstrating that explicitly modeling inter-industry lead–lag relationships is essential for strong predictive accuracy. 3) Removing the skip-connection module in Section 3.6 also reduces performance, confirming its crucial role in facilitating information transmission and gradient flow, thereby improving the model’s feature extraction capabilities. 4) When the Predictor relies solely on the initial node embedding from Equation 27 without integrating the updated features from Equation 28, the model performance undergoes a significant degradation. This is expected given the inherent complexity of stock forecasting, where primitive encodings are insufficient to capture intricate market dynamics. Such a sharp decline underscores the necessity and effectiveness of our proposed hypergraph framework.
@@ 4.4 Efficiency Analysis
In order to ensure good applicability of the model in practical applications, the runtime and memory usage of the model are crucial. We conducted an efficiency comparison analysis of Hermes and other hypergraph-based methods, STHAN-SR, ESTIMATE, and ASHyper—see Table 3.
@@ 5 Conclusions
Stock time series forecasting plays a pivotal role in the modern economy, serving as a key tool for investors, financial institutions, and policymakers to make accurate decisions. However, existing forecasting methods often fall short in fully considering the lead-lag between industries and multi-scale information, resulting in limited predictive performance. To address these challenges, this study introduces the Hermes framework. By ingeniously fusing multi-scale analysis with lead-lag modules within a hypergraph network, the Hermes framework significantly boosts the performance of stock forecasting. Specifically, the hyperedge-based moving aggregation module incorporates a sliding window and utilizes dynamic temporal aggregation operations to consider lead-lag dependencies among industries. Meanwhile, the hyperedge-based multi-scale fusion module decomposes raw data at multiple scales, and then employs cross-scale, edge-to-edge message passing to integrate information from different scales while preserving the consistency of each scale. Experimental results on multiple stock datasets show that Hermes outperforms existing state-of-the-art methods.
@@ A.1.1 Datasets
To validate the performance of Hermes, we use three publicly available stock market datasets, with the dataset statistics provided in Table 4. These datasets provide a diverse set of stock time series, allowing us to comprehensively assess the model’s effectiveness across different financial scenarios. We evaluate our approach using three real-world datasets from the US stock market. These datasets contain relatively complete stock-industry relationships or Wiki-based company relations. The NASDAQ and NYSE datasets Feng et al. [2019b] include EOD data from 01/02/2013 to 12/08/2017, filtered from the respective markets. Abnormal patterns and penny stocks were removed, while preserving their representative properties, with NASDAQ being more volatile and NYSE more stable. The S&P500 dataset Huynh et al. [2023] gathers historical price data and industry information for the companies listed in the S&P500 index from the Yahoo Finance database, covering the period from 01/04/2016 to 05/25/2022.
@@
A.1.2 Metrics
Previous studies use different evaluation metrics, making it challenging to perform a comprehensive comparison of various methods. To provide a thorough assessment of the methods performance, we use four of the most commonly applied and reliable metrics: IC and ICIR, which are rank-based evaluation metrics; and SR, which is return-based.
@@ A.1.2 Metrics
  • Sharpe Ratio (SR): Sharpe ratio Sharpe [1994] measures the profitability of the investment method and take into account risk.
@@ A.1.3 Baselines
We conduct a comparative analysis between our proposed Hermes and a range of baseline methods, which include prominent approaches in both machine learning and deep learning methods for STSF. Given that existing research has demonstrated the limited effectiveness of traditional statistical methods Fan and Shen [2024]; Li et al. [2024], we exclude them from our comparative evaluation to focus on more competitive approaches.
@@
  • LSTM Hochreiter and Schmidhuber [1997] refers to the standard LSTM model, which operates on sequential data, including closing prices and moving averages over 5, 10, 20, and 30 days, to generate a sequential embedding. A fully connected (FC) layer is then employed to predict the return ratio.
@@
  • ALSTM Feng et al. [2019a] integrates adversarial training and stochastic simulation into an enhanced LSTM model, allowing it to more effectively capture market dynamics.
@@
  • RGCN Li et al. [2021] introduces a model that captures both positive and negative correlations among stocks using a correlation matrix computed from historical market data. The LSTM mechanism in RGCN layers helps mitigate over-smoothing issues when predicting overnight stock price movements.
@@
  • GAT Veličković et al. [2017] employs Graph Attention Networks (GAT) to aggregate stock embeddings, which are encoded by a GRU, on the stock graph. By combining the advantages of Graph Neural Networks (GNN) and attention mechanisms, it enhances performance in large-scale graph settings.
@@
  • RSR Feng et al. [2019b] introduces a novel neural network-based framework called Relational Stock Ranking, which addresses the stock forecasting problem in a learning-to-rank approach. Additionally, it proposes a new component in neural network modeling, called Temporal Graph Convolution, which is designed to explicitly capture the domain knowledge of stock relationships in a time-sensitive manner.
@@
  • STHAN-SR Sawhney et al. [2021] introduces a novel Spatio-Temporal Hypergraph Attention Network that models inter-stock relationships of different types and strengths as a hypergraph for stock ranking. It combines temporal Hawkes attention with spatial hypergraph convolutions through hypergraph attention to capture both the correlations in the movements of related stocks and the temporal evolution of their historical features.
@@
  • ESTIMATE Huynh et al. [2023] integrates temporal generative filters within a memory-based shared parameter LSTM network, enhancing the model’s ability to learn temporal patterns for each stock. Furthermore, it introduces attention hypergraph convolutional layers that utilize the wavelet basis—a convolution approach that leverages the polynomial wavelet basis to streamline message passing and focus on localized convolutions.
@@
  • Linear Zeng et al. [2023] employs a straightforward approach by utilizing only fully connected layers to predict the final price. This method does not incorporate any advanced architectures or temporal dependencies, relying solely on basic linear transformations to generate the forecasting. Despite its simplicity, it serves as a baseline model for comparison with more complex models.
@@
  • StockMixer Fan and Shen [2024] introduces a lightweight yet effective MLP-based architecture for stock price forecasting. The model incorporates three key components—indicator mixing, time mixing, and stock mixing—which work together to capture the intricate correlations within the stock data.
  • DLinear Zeng et al. [2023] employs a straightforward linear regression approach that challenges complex Transformer-based architectures. It decomposes the time series and directly regresses historical data for future prediction using a weighted sum operation, demonstrating that simple linear formulations can effectively capture long-term temporal trends.
  • iTransformer Liu et al. [2024] applies attention and feed-forward networks on inverted dimensions to better capture multivariate correlations. Specifically, it embeds the time series of each variable into variate tokens to model dependencies among variables via attention, while utilizing feed-forward networks to learn nonlinear representations for each token.
  • TimeMixer Wang et al. [2024] utilizes a fully MLP-based architecture designed to leverage disentangled multiscale time series. It incorporates Past-Decomposable-Mixing (PDM) blocks to mix seasonal and trend components across different scales, and Future-Multipredictor-Mixing (FMM) blocks to aggregate predictions, thereby capturing both microscopic and macroscopic information effectively.
  • PatchTST Nie et al. [2023] presents an efficient Transformer-based design centered on patching and channel independence. It segments time series into subseries-level patches to serve as input tokens, allowing the model to attend to longer histories with reduced computational costs while retaining local semantic information through independent channel processing.
  • ASHyper Shang et al. [2024] proposes an Adaptive Multi-Scale Hypergraph Transformer to model group-wise interactions in time series. It features an adaptive hypergraph learning module and a multi-scale interaction module to capture comprehensive patterns. Additionally, it employs node and hyperedge constraints to cluster nodes with similar semantics and differentiate temporal variations.
  • DUET Qiu et al. [2025b] introduces a dual clustering framework that operates on both temporal and channel dimensions to address pattern heterogeneity. It employs a Temporal Clustering Module (TCM) to group time series into fine-grained distributions and a Channel Clustering Module (CCM) to capture inter-channel relationships in the frequency domain, effectively mitigating noise and modeling complex dependencies.
@@ B.1 Traditional Stock Time Series Forecasting
Early stock forecasting methods rely mainly on statistical learning techniques to capture patterns and relationships in time series Ye et al. [2016]; Ariyo et al. [2014]; Gupta and Dhingra [2012]; Elton et al. [2009]. One of the most widely used techniques is the Autoregressive Integrated Moving Average (ARIMA) Ariyo et al. [2014] model, which stabilizes data by combining Autoregressive components, Moving Average, and differencing, thereby modeling time series. Another commonly used traditional method is the Hidden Markov Model (HMM) Gupta and Dhingra [2012], which is used to describe unobservable states in the market and can detect changes in market conditions, making it suitable for analyzing market transitions. Although these methods provide predictive capabilities, they exhibit significant limitations when dealing with nonlinear, dynamic changes, and sudden events in financial markets. As a market environment becomes more complex, traditional statistical methods increasingly fail to provide sufficient predictive accuracy and robustness.
With the rapid development in machine learning, machine learning methods for stock time series forecasting have emerged Alkhatib et al. [2013]; Nugroho et al. [2014]; Kim [2003]. For example, Support Vector Machines (SVMs) Kim [2003] map data to high-dimensional spaces using kernel functions and find optimal separating hyperplanes in those spaces, which allows them to capture complex relationships and dynamics in financial markets. However, SVMs are highly sensitive to parameter selection and often struggles to capture long-term dependencies. Overall, although machine learning methods show considerable potential in stock forecasting, they still rely on manual feature engineering and model design. Moreover, these traditional stock forecasting methods have not delved into the exploration of lead-lag relationships and multi-scale modeling.
@@ B.2 Deep Learning Stock Time Series Forecasting
With the rise of deep learning, notable progress has occurred in stock forecasting. This is mainly because deep neural networks can model complex nonlinear patterns. Some studies use recurrent neural networks (RNNs) Nelson et al. [2017]; Qin et al. [2017]; Akita et al. [2016]; Rahman et al. [2019] to capture the historical development in individual stock prices and predict their short-term trends. For example, Long Short-Term Memory (LSTM) networks combined with historical stock data and text information are used to predict stock prices Akita et al. [2016]. In addition, due to the significant progress of Multi-Layer Perceptrons (MLPs) architecture in the general time series field, the DLinear Zeng et al. [2023], CycleNet Lin et al. [2025], and SparseTSF Lin et al. [2024] show the benefits of this simple architecture for time series forecasting. Increasingly many proposals target stock forecasting performance by improving the MLPs architecture. For example, a series of studies Fan and Shen [2024]; Lazcano et al. [2024]; Tashakkori et al. [2024] enhance the learning ability of simple MLPs structures by utilizing the MLP-Mixer backbone, thereby improving forecasting accuracy. Next, the Transformer model has become increasingly popular due to the ability of its self-attention mechanism to capture long-term dependencies. For example, Master Li et al. [2024] proposes a novel stock transformer for stock price forecasting to effectively capture stock correlation.
Some deep learning models have begun to explore lead-lag relationships and multi-scale modeling. For instance, StockMixer Fan and Shen [2024] engages in multi-scale modeling by segmenting the original time series into subsequence-level patches and mixing features at different scales. However, this approach lacks effective integration of information across scales; it merely concatenates information from different scales without thorough fusion. On the other hand, ESTIMATE Huynh et al. [2023] takes into account lead-lag relationships through data-driven detection of leading and lagging time series clusters. Nonetheless, this method falls short in capturing fine-grained lead-lag dynamics, as clustering can only reveal global lead-lag relationships. Distinctive from the aforementioned approaches, Hermes employs a sliding window technique and utilizes dynamic temporal aggregation operations to consider fine-grained lead-lag dependencies among stock industries. Moreover, it features a multi-scale fusion module, which utilizes a cross-scale, edge-to-edge message-passing to integrate information from different scales while maintaining the consistency of each scale.